# Should AI Optimize Your Code? A Comparative Study of Current Large Language Models Versus Classical Optimizing Compilers

### v1

@@ Abstract.
In the contemporary landscape of computer architecture, the demand for efficient parallel programming persists, needing robust optimization techniques. Traditional optimizing compilers have historically been pivotal in this endeavor, adapting to the evolving complexities of modern software systems. The emergence of Large Language Models (LLMs) raises intriguing questions about the potential for AI-driven approaches to revolutionize code optimization methodologies.
This paper presents a comparative analysis between two state-of-the-art Large Language Models, GPT-4.0 and CodeLlama-70B, and traditional optimizing compilers, assessing their respective abilities and limitations in optimizing code for maximum efficiency. Additionally, we introduce a benchmark suite of challenging optimization patterns and an automatic mechanism for evaluating performance and correctness of the code generated by such tools. We used two different prompting methodologies to assess the performance of the LLMs – Chain of Thought (CoT) and Instruction Prompting (IP). We then compared these results with three traditional optimizing compilers, CETUS, PLUTO and ROSE, across a range of real-world use cases.
A key finding is that while LLMs have the potential to outperform current optimizing compilers, they often generate incorrect code on large code sizes, calling for automated verification methods. Our extensive evaluation across 3 different benchmarks suites shows CodeLlama-70B as the superior optimizer among the two LLMs, capable of achieving speedups of up to 2.1x. Additionally, CETUS is the best among the optimizing compilers, achieving a maximum speedup of 1.9x. We also found no significant difference between the two prompting methods: Chain of Thought (Cot) and Instructing prompting (IP).
Automatic parallelization, Large Language Models, Compilers
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: March 01–05, 2025; Las Vegas, NV; isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Parallel programming languagesccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Software verification and validation
@@ 1. Introduction
The need for developing and optimizing parallel programs is paramount (Leiserson et al., 2020). Optimizing compilers have long been one of the pillars of creating parallel code, evolving continuously to meet the growing complexity of modern programs. However, despite all advancements, current compilers still do not achieve the necessary performance to be true alternatives to manual parallelization (Bhosale et al., 2022).
The dawn of Large Language Models (LLMs), such as GPT and CodeLlama raises a fundamental question: Can AI-driven models revolutionize the way we approach code optimization? These models paint an appealing horizon for code optimization and parallel code generation  (Pornmaneerattanatri et al., 2024). Among several LLMs, we chose to use GPT4.0 from OpenAI  (OpenAI, [n. d.]) and CodeLlama-70B  (META, [n. d.]) from META, as these represent the state-of-the-art and thus offer the promise of efficient code generation  (Liu et al., 2023).
@@
In our evaluation, we are specifically interested in code patterns that challenge current automatic optimization tools and in the ability of LLM-based AOTs to overcome these challenges. To this end, we have developed a benchmark suite of such patterns. Next to a comprehensive evaluation, the suite also allows us to identify areas for improvement in parallel code optimization.
@@
  • We introduce an automated mechanism for validating the correctness and performance of AOT-generated code.
  • We present a Parallel Computing Challenge Benchmark suite (PCB) v1.0, comprising 20 use cases for evaluating the capabilities of AOTs.
This paper is organized as follows: Section  2 explores strengths and limitations of our chosen optimizing compilers and LLMs. Section  3 describes our environment for evaluating automatic optimization tools and presents the Parallel Computing Challenge Benchmark suite (PCB). Section  4 evaluates the capabilities of our LLMs and optimizing compilers on challenging code patterns, followed by the discussion of related work in Section 5 and conclusions  6.
@@ 2. LLMs vs Traditional Compilers, Strengths and Limitations
Current optimizing compilers exhibit a mix of challenges and opportunities (Monsifrot et al., 2002; Zhao et al., 2005; Li et al., 2014). They include limitations of static analysis, hampering the identification of optimization opportunities, and the complexity of deciding on best optimization patterns for maximum program performance (Barakhshan and Eigenmann, 2023b). It is in this context that the power of LLMs emerges, offering new potential ways to approach program optimization.
Large Language Models have garnered significant attention in recent years owing to their statistical understanding of programs, offering suggestions for code generation, and facilitating software creation (Li et al., 2023) (Li et al., 2022) (Achiam et al., 2023) (Chen et al., 2021). While LLM-generated software has demonstrated success in many applications, recent studies have highlighted the limitations in capturing crucial aspects of linguistic meaning and understanding semantic properties of the user’s prompt (Asher et al., 2023).
Despite these limitations, recent studies indicate that Large Language Models excel in programming tasks such as code generation, issue fixing, and code completion (Valero-Lara et al., 2023). This proficiency is largely attributed to the vast amount of data available on platforms such as GitHub, enabling effective training. However, even though LLMs can generate effective code, the lack of correctness guarantees make them unsuitable for many code optimization approaches (Shypula et al., 2023; Kadosh et al., 2023; Chen et al., 2023; Nichols et al., 2023).
In pursuit of the question asked in the Introduction Can AI-driven models revolutionize the way we approach code optimization?, this paper investigates in what situations Large Language Models perform well and where not. To this end, we have developed an environment for comparing the optimization capabilities of LLMs with those of classical optimizing compilers, including the creation of an automated code verification mechanism and a benchmark suite of code challenge patterns. The next section describes this environment.
@@ 3. An Environment for Evaluating Automatic Optimization Tools
Recall that the lack of a methods for verifying the correctness of LLM-generated code make it difficult to compare such tools with other AOTs. Section  3.1 introduces a novel method and environment called PCAOT (Performance and Correctness Evaluation of Automatic Optimization Tools) that overcomes this limitation. Additionally, we have created the Parallel Challenge Patterns Benchmark (PCB) suite v1.0 with 20 distinct use cases. The suite and the challenge patterns are discussed in Section  3.2.
@@ 3.1. Evaluation and Verification
Our approach entails selecting two Large Language Models (LLMs), GPT-4.0 and CodeLlama-70B, and evaluating their code optimization capabilities. The results are then compared with three different optimizing compilers, CETUS, PLUTO and ROSE. For engaging the LLMs, we use two prompting strategies, descibed next.
@@ 3.1.1. Prompting Strategies
We explored two different prompts associated with utilizing Large Language Models for High-Performance Computing optimization tasks. Initially, we employed instruction prompting (IP)  (Mishra et al., 2021)  (Gupta et al., 2022). Additionally, we employed Chain of Thought (CoT) within these Large Language Models. Recent studies indicate that Chain of Thought prompting enhances performance across a spectrum of arithmetic, commonsense, and symbolic reasoning tasks, making it well suited for assessing domain-specific knowledge such as Parallel Computing  (Wei et al., 2022).
@@
  • Instructing (IP) - ”Given the program below, improve its performance using OpenMP”.
  • CoT - ”Given the C program below, think of a way how to optimize its performance using OpenMP”
The distinction between these two types of prompts lies in their approach to guiding the model’s reasoning process. Instructing prompts provide a direct and clear directive, resulting in the immediate production of an optimized program version. In contrast, Chain of Thought prompts enable the model to break down the task into multiple steps, allowing for iterative refinement before generating an optimized program version. Instances of guidance the model in this process are: ”Think of a way…”, ”propose…”, ”give a potential strategy…”  (Shypula et al., 2023).
For the optimizing compilers CETUS, PLUTO, and ROSE, we used the default set of parameters, representing the known, safe optimization methods. Note that, while experimenting with optimization options could potentially yield improved performance, doing so is beyond the scope of this paper. Our environment for evaluating the five AOTs comprises three phases: Preparation, Optimization, and Validation. Figure 1 illustrates this architecture.
@@ 3.1.2. Preparation Phase
LLMs have shown to have limitations to process programs with large code sizes  (OpenAI, [n. d.])  (META, [n. d.]). Hence, our LLM evaluation focuses on section-level optimizations. The preparation phase identifies the sections within the source code to experiment with. Each such section is manually enclosed with #pragma experimental section start and  #pragma experimental section end, as illustrated in Figure 2. The impact of code size on the performance of the LLMs will be discussed in Section 4.2.
@@
  10:                rhs[k][j][i][m]=forcing[k][j][i][m];
  11:            }
  12:        }
  13:    }
  14:}
  15: #pragma experimental end
Figure 2. Experimental section instrumentation. Example is from subroutine Compute_rhs of the NPB Application CG.
For our experiments we made used of a tool called CaRV (Capture, Replay, and Validate), which enables users to experiment with sections of large applications. It facilitates the comparison of individual program sections before and after optimizations, assessing their efficiency and accuracy  (Barakhshan and Eigenmann, 2023a). CaRV does so using a selective checkpointing method, based on the identified program states at the beginning and end of an experimental section.
The preparation phase generates a new program version that is instrumented in a way that allows the optimized program section to be run independently and compared against the captured end state of the original program. Both the original and instrumented program versions are passed on to the optimization phase.
@@ 3.1.3. Optimization Phase
This phase sends the experimental section with the prompts to the two LLMs to obtain the optimized versions. The resulting code is then inserted into the instrumented program. The original code is also sent to the three compilers, CETUS, PLUTO and ROSE. This process generates a total of 16 program versions, for the five AOTs (two LLMS, two different prompts, three requests per prompt, three optimizing compilers plus the serial program) The next phase executes and validates these versions.
@@ 3.1.4. Validation Phase
The Validation phase executes each version of the experimental section individually. The LLM-generated versions are then compared with the original, serial program versions, using the CaRV checkpoints. This process verifies the values of all relevant, ”live” variables at the end of the LLM-optimized program section against the corresponding values in the original program. The execution time is measured as well and compared to that of the original program section.
@@
This new Benchmark suite comprises 20 distinct use cases, each presenting one of six challenging scenarios encountered by optimizing compilers  (Barakhshan and Eigenmann, 2023b). A number of patterns are included in the PCB suite, testing the optimizers in their ability to
  • parallelize outermost loops (PO),
  • parallelize loops with function calls (PF),
  • form parallel regions enclosing multiple parallel loops (PR),
  • avoiding load imbalance through dynamic scheduling (DS),
  • parallelizing array reductions (PA),
  • eliminate barrier synchronizations using NOWAIT clause (NW).
@@
Given the challenges traditional optimizing compilers face in handling these patterns (Barakhshan and Eigenmann, 2023b)  (Prema et al., 2017), an important question is: Can LLMs identify these patterns and perform the requisite optimizations? The following paragraphs explain the involved program patterns:
Parallelizing Outermost Loops
This pattern involves identifying and parallelizing the outermost loops in a program section, which often represent significant computational tasks. Parallelizing these loops can improve performance by distributing the number of iterations across multiple processors or threads. Running the outer loop in parallel typically results in lower overhead compared to parallelizing the inner loop, as inner parallelism increases the number of times of starting and ending parallel activities. This pattern poses a challenge because real-world scenarios often contain loops with complex expressions that prove too difficult for compilers’ static analysis capabilities. Automatic optimizers often success in parallelizing the inner loops only. Figure  3 illustrates this situation.
    1: int sum
    2: for (int i = 0; i < size; ++i) {
    3:   for (int j = 0; j < size; ++j) {
    4:      for (int k = 0; k < size; ++k) {
    5:       #pragma omp parallel for
    6:       for (int l = 0; l < size; ++l) {
    7:         int position =i*size * size * size + j * size ;
    8:         int index = position  * size + k * size + l;
    9:         sum += data[indices[index]];
    10:        }
    11:     }
    12:   }
    13: }
Figure 3. Parallelizing the outermost loop pattern in an example loop from the PCB v1.0. Compilers tend to parallelize the inner-most loop, due to the irregular data accesses in code line #9. Doing so incurs the overhead of starting the parallel region $size^{3}$ Parallelizing the outermost loop would eliminate this overhead.
Parallelizing Loops with Function Calls
Parallelizing loops with function calls presents a significant challenge for optimizing compilers, as many optimization techniques do not operate interprocedurally. A special situation is the presence of calls to side-effect-free functions (functions that only modify their parameters), which only require the optimizer to determine the independence of the variables passed as function parameters. Loops with such functions can be parallelized. Figure 4 illustrates this situation.
    1: int size = 1000;
    2: int result[size][size];
    3: for (int i = 0; i < size; i++) {
    4:    for (int j = 0; j < size; j++) {
    5:        for (k = 0; k < size; k++){
    6:            result[i][j][k] = compute(i, j, k);
    8:            }
    9:    }
    10: }
    13: /* Side effect free fuction call*/
    14: int compute(int x, int y, int k) {
    15:    int result = 0;
    16:       for (int i = 0; i < x; i++) {
    17:        result += (x + y + k) * i;
    20:        return result;
    21:    }
    22: }
Figure 4. Parallelizing Loops with Function Calls pattern in an example loop from the PCB v1.0. Function compute is side effect free. The loop in line #3 can be parallelized.
Parallel Regions Enclosing Multiple Parallel Loops
Current optimizing compilers are limited to optimizing loops individually and cannot apply optimizations across a set of loops within a parallel region. By enclosing multiple parallel loops within a single parallel region, the overhead associated with starting and ending each parallel loop separately is reduced. Figure  5 illustrates this pattern.
    1: int array[1000];
    2: int result[1000];
    3:  // Initialize array
    4:    for (int i = 0; i < NUM_ELEMENTS; i++) {
    5:        array[i] = i;
    6:    }
    7:
    8:   #pragma omp parallel
    9: {
    10:    #pragma omp parallel for
    11:    for (int i = 0; i < NUM_ELEMENTS; i++) {
    12:        result[i] = array[i] * 2;
    13:    }
    14:
    15:    #pragma omp parallel for
    16:    for (int i = 0; i < NUM_ELEMENTS; i++) {
    17:       result[i] *= result[i];
    18:    }
    19:
    20:    #pragma omp parallel for
    21:    for (int i = 0; i < NUM_ELEMENTS; i++) {
    22:       result[i] += 10;
    23:    }
    24: }
Figure 5. Parallel Regions Enclosing Multiple Parallel Loops pattern in an example from the PCB v1.0. The parallel region reduces the Fork/join overhead caused by parallelizing each loop individually.
Avoiding Load Imbalance Through Dynamic Scheduling
Load imbalance occurs when the computational workload is unevenly distributed among the available processing units. Load imbalance is a significant source of performance degradation and limited scalability in parallel computing. Dynamic scheduling in OpenMP can mitigate this effect by assigning computational tasks to processing units at runtime based on their load and availability. Figure 6 illustrates this situation by showcasing a triangular loop. The default OpenMP scheduler (static) will divide the iterations of the parallel loop evenly, causing load imbalance.
    1: #pragma omp parallel for schedule(dynamic, 1)
    2: //Every thread will process one iteration of the loop,
    3: // and it will be used as it is available
    4:  for (int i = 0; i < N; i++) {
    5:    for (int j = 0; j < N - i; j++) {
    6:      // With static scheduling: Thread1->N-i,
    7:        // Thread2->N-i-1...
    8:      // ThreadN->1
    9:       array[i][j] = i + j;
   10:    }
   11:  }
Figure 6. Avoiding Load Imbalance Through Dynamic Scheduling pattern in an example from the PCB v1.0. Inserting the OpenMP clause for dynamic scheduling resolves the load imbalance.
Parallelizing Array Reductions
Array reductions perform operations, such as sum, product, min, or max, on elements of an array. Parallelizing these operations requires careful consideration of data dependencies and synchronization to ensure correct results while maximizing parallel execution. Algorithms to perform reductions in parallel are well known (Pottenger and Eigenmann, 1995) but applying them correctly and beneficially when the result is an array itself can be non trivial (Prema et al., 2017). Figure 7 illustrates this situation.
    1:  #pragma omp parallel for reduction(+:result)
    2:  for (i = 0; i < ARRAY_SIZE; i++) {
    3:       for (j = 0; j < INNER_SIZE; j++) {
    5:          result[j] += array[i][j];
    6:       }
    7:    }
    8:
Figure 7. Parallelizing Array Reduction pattern in an example from the PCB v1.0. Most compilers can recognize scalar reductions. The analysis is more difficult, if the reduction variable is an array, as result[] in this example.
NOWAIT – Eliminating Barrier Synchronizations
Barriers are essential for maintaining correctness and synchronization in parallel programs but introduce overhead when some threads complete their work before others. NOWAIT constructs can be used to eliminate barrier synchronizations and improve parallel performance by allowing threads to continue executing. AOTs need to determine the situations where doing so is legal. Figure 8 illustrates this pattern.
    1: #pragma omp parallel
    2: {
    3:    double local_sum = 0.0;
    4:    for (int i = 0; i < ARRAY_SIZE; i++) {
    5:        for (int j = 0; j < INNER_SIZE; j++) {
    6:             for (int k = 0; k < INNER_SIZE; k++) {
    7:                result[i][j][k]+=matrix_A[i][j][k]
    8:               * matrix_B[i][j][k];
    9:              }
    10:        }
    11:   }
    12:       #pragma omp for nowait
    13:       for (i = 0; i < ARRAY_SIZE; i++) {
    14:         for (j = 0; j < INNER_SIZE; j++) {
    15:             for (k = 0; k < INNER_SIZE; k++) {
    16:                 local_sum += array[i][j][k];
    17:             }
    18:         }
    19:     }
    20:    #pragma omp atomic
    21:     sum += local_sum;
    22:   }
Figure 8. Eliminating Barrier Synchronizations pattern in an example from the PCB v1.0. An OpenMP NOWAIT clause is inserted to eliminate the barrier synchronization and improve parallel performance.

### v2

# Should AI Optimize Your Code? A Comparative Study of Classical Optimizing Compilers Versus Current Large Language Models

@@ Abstract.
Traditional optimizing compilers have played an important role in adapting to the growing complexity of modern software systems. The need for efficient parallel programming in current architectures requires strong optimization techniques. The beginning of Large Language Models (LLMs) raises intriguing questions about the potential of these AI approaches to revolutionize code optimization methodologies. This work aims to answer an essential question for the compiler community: ”Can AI-driven models revolutionize the way we approach code optimization?”.
To address this question, we present a comparative analysis between three classical optimizing compilers and two recent large language models, evaluating their respective abilities and limitations in optimizing code for maximum efficiency. In addition, we introduce a benchmark suite of challenging optimization patterns and an automatic mechanism for evaluating the performance and correctness of the code generated by LLMs. We used three different prompting strategies to evaluate the performance of the LLMs – Simple Instruction (IP), Detailed Instruction Prompting (DIP), and Chain of Thought (CoT).
A key finding is that while LLMs have the potential to outperform current optimizing compilers, they often generate incorrect code on large code sizes, calling for automated verification methods. In addition, expressing a compiler strategy as part of the LLM’s prompt substantially improves its overall performance. Our evaluation across three benchmark suites shows CodeLlama-70B as the superior LLM, capable of achieving speedups of up to x1.75. Additionally, CETUS is the best among the current optimizing compilers, achieving a maximum speedup of 1.67x. We also found substantial differences among the three prompting strategies.
Automatic parallelization, Compilers, Large Language Models
journalyear: 2025copyright: rightsretainedconference: SupercomputingAsia 2025; March 10–13, 2025; Singapore, Singaporebooktitle: SupercomputingAsia 2025 (SCA ’25), March 10–13, 2025, Singapore, Singaporedoi: 10.1145/3718350.3718357isbn: 979-8-4007-1250-0/25/03ccs: Computing methodologies Parallel programming languagesccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Software verification and validation [^1] Both authors contributed equally to this research. Accepted at SupercomputingAsia (SCA’25), March 10-13,2025
@@ 1. Introduction
The need for developing and optimizing parallel programs is more crucial than ever (Leiserson et al., 2020). Optimizing compilers have long been one of the pillars of creating parallel code, evolving continuously to meet the growing complexity of modern programs and architectures. However, despite all advancements, current compilers still do not achieve the necessary performance to be true alternatives to manual parallelization (Bhosale et al., 2022).
The dawn of Large Language Models (LLMs), such as GPT and CodeLlama raises a fundamental question: Can AI-driven models revolutionize the way we approach code optimization? These models paint an appealing horizon for code optimization and parallel code generation  (Pornmaneerattanatri et al., 2024). Among several LLMs, we chose to explore GPT4.0 from OpenAI  (OpenAI, [n. d.]) and CodeLlama-70B  (META, [n. d.]) from META, as these represent the state-of-the-art and thus offer the promise of efficient code generation  (Liu et al., 2023).
@@
In our evaluation, we are specifically interested in code patterns that challenge current automatic optimization tools and in the ability of LLM-based AOTs to overcome these challenges. To this end, we have developed a benchmark suite of such patterns. Next to a comprehensive evaluation, the suite is also available to identify areas for improvement in parallel code optimization.
@@
  • We introduce an automated mechanism for validating the correctness and performance of AOT-generated code called PCAOT (Performance and Correctness Evaluation of Automatic Optimization Tools).
  • We present a Parallel Computing Challenge Benchmark suite (PCB) v1.0 (Miguel Rosas, Miguel Torres, [n. d.]), comprising 28 parallel and non-parallel use cases for evaluating the capabilities of AOTs.
This paper is organized as follows: Section  2 explores strengths and limitations of our chosen optimizing compilers and LLMs. Section  3 describes our environment for evaluating automatic optimization tools and presents the Parallel Computing Challenge Benchmark suite (PCB). Section  4 evaluates the capabilities of our LLMs and the optimizing compilers on challenging code patterns for both parallel and non-parallel scenarios, followed by the discussion of related work in Section 5 and conclusions in Section  6.
@@ 2. LLMs vs Traditional Compilers, Strengths and Limitations
Current optimizing compilers exhibit a mix of challenges and opportunities (Monsifrot et al., 2002; Zhao et al., 2005; Li et al., 2014). Chief among them are limitations of static analysis, hampering the identification of optimization opportunities, and the complexity of deciding on best optimization patterns for maximum program performance (Barakhshan and Eigenmann, 2023b). It is in this context that the power of LLMs emerges, offering new potential ways to approach program optimization.
Large Language Models have captured significant attention in recent years owing to their statistical understanding of programs, offering suggestions for code generation, and facilitating software creation (Li et al., 2023), (Li et al., 2022), (Achiam et al., 2023), (Chen et al., 2021). While LLM-generated software has demonstrated success in many applications, recent studies have highlighted the limitations in capturing crucial aspects of linguistic meaning and understanding semantic properties of the user’s prompt (Asher et al., 2023).
Regardless of these limitations, recent studies suggest that Large Language Models perform remarkably well in different programming tasks, such as code generation, bug fixing, and code completion (Valero-Lara et al., 2023). This expertise is largely attributed to the amount of data available on platforms such as GitHub, which provides an effective training. However, even though LLMs can generate effective code, the lack of correctness guarantees make them unsuitable for many code optimization approaches (Shypula et al., 2023; Kadosh et al., 2023; Chen et al., 2023; Nichols et al., 2023).
@@ 3. An Environment for Evaluating Automatic Optimization Tools
Recall that the lack of methods for verifying the correctness of code generated by LLMs makes it difficult to compare such tools with other AOTs. Section  3.1 introduces a novel method and environment called PCAOT (Performance and Correctness Evaluation of Automatic Optimization Tools) that overcomes this limitation. Additionally, we have created the Parallel Challenge Patterns Benchmark (PCB) suite v1.0 with 28 parallel and non-parallel use cases. The suite and the challenge patterns are discussed in Section  3.2.
@@ 3.1. Evaluation and Verification
Our approach entails selecting two Large Language Models (LLMs), GPT-4.0 and CodeLlama-70B, and evaluating their code optimization capabilities. The results are then compared with three different optimizing compilers, CETUS, PLUTO, and ROSE. For engaging the LLMs, we use three prompting strategies, described next.
@@ 3.1.1. Prompting Strategies
We explored three different prompts associated with utilizing Large Language Models for High-Performance Computing optimization tasks. Initially, we employed a simple instruction prompt (IP)  (Mishra et al., 2021)  (Gupta et al., 2022). Additionally, we transformed the original simple instruction prompt into a more detailed one (DIP). Finally, we employed a Chain-of-Thought prompt (CoT). Recent studies indicate that Chain-of-Thought prompting enhances performance across a spectrum of arithmetic, commonsense, and symbolic reasoning tasks, making it well suited for assessing domain-specific knowledge  (Wei et al., 2022).
@@
  • Simple Instructing Prompting (IP) - ”Given the program below, improve its performance using OpenMP.”
  • Detailed Instructing Prompting (DIP) - ”Given the C program below, check for read after write and write after read dependencies among iterations, if there are no dependencies among iterations of the outermost loop, parallelize this loop using OpenMP directives. If dependencies are found in the outermost loop but there exist inner loops that can be parallelized without violating data dependencies, then parallelize those inner loops instead.”
  • Chain-of-Thought Prompt - ”¡DIP¿ + As you work through the program, explain each step of your reasoning process to ensure clarity and correctness in your optimization decisions. Think step by step.”
The main difference between the three types of prompts recline in their ability to guide the model’s reasoning process. Instructing prompts provide a more direct and clear directive, resulting in the immediate production of an optimized program version. As opposed to Chain of Thought (CoT) prompts, which enable the model to break down the task into multiple steps, allowing for iterative refinement before generating an optimized program version. Instances of guidance in the model in this process are: ”Explain each step of your reasoning process…”, ”Break down your thought process into individual steps…”, ”Detail the reasoning behind your answer…”  (Shypula et al., 2023).
We chose to add a condition into the prompts DIP and CoT which is ”check for read after write, and write after read dependencies among iterations, If there are no dependencies among iterations of the outermost loop, parallelize this loop using OpenMP directives….”. This guidance consists of two key components: first, analyzing data dependencies to ensure correctness, and second, prioritizing the parallelization of outer loops. These key components represent two of the most fundamental concepts in automatic program parallelization. Additionally, when prompted simply to ”improve its performance”, with IP, LLMs are likely to attempt parallelization regardless of whether the code is actually parallelizable. The DIP and CoT prompts express high level compiler strategies that guide the LLMs to improve their reasoning process.
For the optimizing compilers, we used the default set of parameters, representing the known, safe optimization methods. Note that, while experimenting with optimization options could potentially yield improved performance, doing so is beyond the scope of this study. Our environment for evaluating the five AOTs comprises three phases: Preparation, Optimization, and Validation. Figure 1 illustrates this architecture.
@@ 3.1.2. Preparation Phase
LLMs have shown to have limitations in processing programs with large code sizes  (OpenAI, [n. d.])  (META, [n. d.]). Hence, our LLM evaluation focuses on section-level optimizations. The preparation phase identifies the sections within the source code to experiment with. Each such section is manually enclosed with #pragma experimental section start and  #pragma experimental section stop, as illustrated in Figure 2. The impact of code size on the performance of the LLMs will be discussed in Section 4.2.
@@
  10: #pragma experimental end
Figure 2. Experimental section instrumentation. Example taken from subroutine PO_Function_Version_1 of the PCB suite.
For our experiments, we made use of a tool called CaRV (Capture, Replay, and Validate), which enables users to experiment with sections of large applications. It facilitates the comparison of individual program sections before and after optimizations, assessing their efficiency and accuracy  (Barakhshan and Eigenmann, 2023a). CaRV does so using a selective checkpointing method, based on the identified program states at the beginning and end of an experimental section. The preparation phase generates a new program version that is instrumented in a way that allows the optimized program section to be run independently and compared against the captured end state of the original program. Both the original and instrumented program versions are passed on to the optimization phase.
@@ 3.1.3. Optimization Phase
This phase sends the experimental section with the prompts to the two LLMs to obtain the optimized versions. The resulting code is then inserted into the instrumented program. The original code is also sent to the three compilers, CETUS, PLUTO, and ROSE. This process generates a total of 22 program versions, for the five AOTs (two LLMS, three different prompts, three requests per prompt, three optimizing compilers plus the serial program). The next phase executes and validates these versions.
@@ 3.1.4. Validation Phase
The Validation phase executes each version of the experimental section individually. The LLM-generated versions are then compared with the original, serial program versions, using the CaRV checkpoints. This process verifies the values of all relevant, ”live” variables at the end of the LLM-optimized program section against the output state of the corresponding values in the original program. The execution time is measured as well and compared to that of the original program section.
@@
The PCB Suite was specifically created to address a critical need for unbiased evaluation by ensuring that it has not been exposed to LLMs during training. This new Benchmark suite comprises 28 distinct use cases, each presenting one of six challenging scenarios encountered by optimizing compilers  (Barakhshan and Eigenmann, 2023b). A number of patterns are included in the PCB suite, testing the optimizers in their ability to: parallelize outermost loops (PO), parallelize loops with function calls (PF), form parallel regions enclosing multiple parallel loops (PR), parallelizing array reductions (PA), avoiding load imbalance through dynamic scheduling (DS), and eliminate barrier synchronizations using NOWAIT clause (NW).
@@
These patterns were identified as key challenge areas for classical optimizing compilers  (Barakhshan and Eigenmann, 2023b),  (Prema et al., 2017), an important question is: Can LLMs identify these patterns and perform the requisite optimizations? More detailed explanation of these challenge patterns can be found in prior studies (Barakhshan and Eigenmann, 2023b).